5. Gravetter/Wallnau/Forzano, Essentials Chapter 1- End-of-chapter question 20 For the following set of scores, find the value of each expression: 2 2 (2X)2- E(x + 3)
Answer #1

Answer :

X X2 X + 3
- 2 4 - 2 + 3 = 1
- 2 4 - 2 + 3 = 1
7 49 7 + 3 = 10
- 3 9 - 3 + 3 = 0
- 1 1 - 1 + 3 = 2

\sumX = (-2) + (-2) + 7 + (-3) + (-1) = - 1

Therefore, \sumX = - 1.

(\sumX)2 = ( - 1 )2 = 1

Therefore, (\sumX)2 = 1.

\sumX2 = 4 + 4 + 49 + 9 + 1 = 67

Therefore, \sumX2 = 67.

\sum(X + 3) = 1 + 1 + 10 + 0 + 2 = 14

Therefore, \sum(X + 3) = 14.